# iA: Imperative Learning-based A Search for Path Planning

### v2

@@ V Conclusion
We presented iA*, an innovative search-based path planning framework designed to address critical challenges in existing methodologies. Leveraging imperative learning (IL), iA* integrates neural networks with a differentiable A* algorithm in a unified bilevel optimization framework, effectively eliminating the need for extensive labeled datasets through self-supervised training. This approach mitigates overfitting and significantly enhances generalization to previously unseen environments. Comprehensive experimental evaluations across diverse datasets and realistic simulation scenarios demonstrated iA*’s superior performance compared to both classical and supervised learning-based methods. The effectiveness of our IL-based training approach was further confirmed by ablation studies, underscoring its robustness and adaptability. Future research will explore extending iA* to higher-dimensional planning tasks and integrating dynamic obstacles, further broadening its applicability.

### v4

@@ V Conclusion and Future Work
We present iA*, an innovative search-based path planning framework designed to address critical challenges in existing methodologies. Leveraging imperative learning (IL), iA* integrates neural networks with a differentiable A* algorithm in a unified bilevel optimization framework, effectively eliminating the need for extensive labeled datasets through self-supervised training. This approach mitigates overfitting and significantly enhances generalization to previously unseen environments. Comprehensive experimental evaluations across diverse datasets and realistic simulation scenarios demonstrated iA*’s superior performance compared to both classical and supervised learning-based methods. The effectiveness of our IL-based training approach is further confirmed by ablation studies, underscoring its robustness and adaptability. In future work, we aim to extend iA* to high-dimensional path planning problems, where the neural model may efficiently address the computational explosion of the search space. Furthermore, enhancing iA* with real-time replanning capabilities will broaden its applicability to dynamic environments, where fast adaptation and responsiveness are crucial for reliable performance.